if a plane is going 240 mph then how fast is it traveling relative to the ground?
Answer by Fombitz(32388)

Depends on the winds aloft and the direction of the plane relative to the wind direction.
Typically airplane speeds are given relative to the ground with zero wind.
So of the wind speed is zero, then V=240 mph.
If the plane travels into a 100mph head wind, then V=240-100=140 mph.
If the plane travels into a 100mph tail wind, then V=240+100=340 mph.